how many moles of hydrogen are needed to completely react with 1.23 moles of nitrogen?
1
Expert's answer
2020-02-03T07:45:42-0500

According to the reaction equation:

N2+3H2=2NH3

n(H2) = n(N2) *3 = 1.23 mol * 3 = 3.69 mol

Answer: 3.69 mol